# JointRF: End-to-End Joint Optimization for Dynamic Neural Radiance Field Representation and Compression

### v1

@@ 2.2 NeRF Compression
Recently, there has been limited attention to compressing static NeRF to reduce storage usage. VQRF[16] employs an entropy encoder to compress the static radiance field model, while ECRF[17] maps radiance field features to the frequency domain before applying entropy encoding. However, these methods remain confined to static scenes, and there has been relatively little exploration in compressing dynamic scenes. ReRF[2] is designed for modeling dynamic scenes and utilizes traditional image compression methods for further feature compression. However, due to the separate representation and compression processes in ReRF[2], it lacks end-to-end optimization. Moreover, traditional image encoders are unsuitable for compressing feature grids, resulting in suboptimal compression efficiency.
@@ 3.1 JointRF Representation
To enhance temporal continuity while minimizing redundancy, we propose a highly compact representation method. The overview of our representation method is illustrated in Fig.2. We represent the long sequence of dynamic radiance field over time as multiple continuous groups of feature grids (GOF), which is a collection of successive grids. Each GOF $\mathbf{G}=\{\mathbf{f}_{t}\}_{t=1}^{N}$ begins with a keyframe $\mathbf{f}_{1}=\{\mathbf{B}_{1},\mathbf{C}_{1}\}$, which is a long reference feature grid. Due to the commonality captured by the basis in signal extraction and the short-term similarity of features, each remaining frame in the GOF is represented as a compact residual grid along with a coefficient grid ($\mathbf{f}_{t}=\{\mathbf{R}_{t},\mathbf{C}_{t}\}$, $1<t\leq N$). The residual feature grid $\mathbf{R}_{t}$ indicates the compensation for errors and newly observed regions between the keyframe’s basis grid and the current frame’s basis grid. $\mathbf{R}_{t}$ helps enhance the representation by incorporating residual information and reducing temporal redundancy. Once obtained $\mathbf{R}_{t}$, we can calculate the basis feature grid of the $t$-th frame by adding the residual compensation:$\mathbf{B}_{t}=\mathbf{B}_{1}+\mathbf{R}_{t}$, enabling the recovery of the current radiance field by applying the signal compensation according to Eq.(2).
@@ 4.2 Comparison
To the best of our knowledge, JointRF is a rare method for jointly training dynamic NeRF and their associated compression processes in a unified end-to-end manner. To validate the effectiveness of our approach, we compare with several state-of-the-art methods for dynamic scenes including INGP[3], MERF[4], TiNeuVox[14], ReRF[2] both qualitatively and quantitatively. In Fig.4, we present the visual results of two sequences. As illustrated, our method demonstrates discernible superiority in both the compactness of the model size and the precision of detail rendering not only on per-frame static reconstruction method INGP[3] and MERF[4] but also on dynamic scene reconstruction methods TiNeuVox[14] and ReRF[2].

### v2

@@ 2.2 NeRF Compression
Recently, there has been limited attention to compressing NeRF to reduce storage usage. VQRF[15] employs an entropy encoder to compress the static radiance field model, while ECRF[16] maps radiance field features to the frequency domain before applying entropy encoding. However, these methods remain confined to static scenes, and there has been relatively little exploration in compressing dynamic scenes. ReRF[2] is designed for modeling dynamic scenes and utilizes traditional image compression methods for feature compression. However, due to the separate representation and compression processes, ReRF[2] lacks end-to-end optimization. Moreover, traditional image encoders are unsuitable for compressing feature grids, resulting in suboptimal compression efficiency. Comparing to [17], our approach introduces different representation and compression methods.
@@ 3.1 JointRF Representation
To enhance temporal continuity while minimizing redundancy, we propose a highly compact representation method. The overview of our representation method is illustrated in Fig.2. We represent the long sequence of dynamic radiance field over time as multiple continuous groups of feature grids (GOF), which is a collection of successive grids. Each GOF $\mathbf{G}=\{\mathbf{f}_{t}\}_{t=1}^{N}$ begins with a keyframe $\mathbf{f}_{1}=\{\mathbf{B}_{1},\mathbf{C}_{1}\}$, which is a long reference feature grid. Due to the commonality captured by the basis in signal extraction and the short-term similarity of features, each remaining frame in the GOF is represented as a compact residual grid along with a coefficient grid ($\mathbf{f}_{t}=\{\mathbf{R}_{t},\mathbf{C}_{t}\}$, $1<t\leq N$). The residual feature grid $\mathbf{R}_{t}$ indicates the compensation for errors and newly observed regions between the keyframe’s basis grid and the current frame’s basis grid. $\mathbf{R}_{t}$ helps enhance the representation by incorporating residual information and reducing temporal redundancy. Once obtained $\mathbf{R}_{t}$, we can calculate the basis feature grid of the $t$-th frame by adding the residual compensation:$\mathbf{B}_{t}=\mathbf{B}_{1}+\mathbf{R}_{t}$, enabling the recovery of the current radiance field by applying the signal compensation according to Eq.(2). It is noteworthy that $\mathbf{C}_{t}$ is the coefficient based on $\mathbf{B}_{t}$, and learning the residuals of $\mathbf{C}_{t}$ holds no physical significance.
@@ 4.2 Comparison
To the best of our knowledge, JointRF is a unique approach for jointly training dynamic NeRF and their associated compression processes in a unified end-to-end manner. To validate the effectiveness of our approach, we compare with several state-of-the-art methods for dynamic scenes including INGP[3], MERF[4], TiNeuVox[9], ReRF[2] both qualitatively and quantitatively. In Fig.4, we present the visual results of two sequences. As illustrated, our method demonstrates discernible superiority in both the compactness of the model size and the precision of detail rendering not only on per-frame static reconstruction method INGP[3] and MERF[4] but also on dynamic scene reconstruction methods TiNeuVox[9] and ReRF[2].